# Adapting Differentially Private Synthetic Data to Relational Databases

### v1

@@ Abstract
Existing differentially private (DP) synthetic data generation mechanisms typically assume a single-source table. In practice, data is often distributed across multiple tables with relationships across tables. In this paper, we introduce the first-of-its-kind algorithm that can be combined with any existing DP mechanisms to generate synthetic relational databases. Our algorithm iteratively refines the relationship between individual synthetic tables to minimize their approximation errors in terms of low-order marginal distributions while maintaining referential integrity. Finally, we provide both DP and theoretical utility guarantees for our algorithm.
@@ 1 Introduction
Relational databases play a pivotal role in modern information systems and business operations due to their efficiency in managing structured data [39]. According to a Kaggle survey [23], 65.5% of users worked extensively with relational data. Additionally, the majority of leading database management systems (e.g., MySQL and Oracle) are built on relational database principles [35]. These systems organize data into multiple tables, each representing a specific entity, and the relationships between tables delineate the connections between these entities. However, the widespread use of relational databases also carries a significant risk of privacy leakage. For example, if a single table suffers from a privacy breach, all other tables containing sensitive information can be exposed, as they are interconnected and share relationships. Moreover, deleting data in a relational database can be complex, and incomplete deletion can leave traces of data, which can potentially be accessed and reconstructed by attackers.
Today, differential privacy (DP) stands as the de facto standard for privacy protection. There is a growing body of research focused on applying DP to generate private synthetic data [see, e.g., 36, 31, 27, 41]. This effort enables data curators to share (synthetic) data while ensuring the privacy of individuals’ personal information within the original dataset. In return, they can harness advanced machine learning techniques employed by end-users trained on the synthetic data. Numerous studies have provided evidence that state-of-the-art (SOTA) DP synthetic data effectively preserves both the statistical properties of the original data and the high performance of downstream predictive models trained on these synthetic datasets when deployed on the original data [45, 51]. However, all existing works assume a single-source database (with a few exceptions discussed in Related Work). It motivates the central question we aim to tackle in this paper:
@@
The conventional approach—flattening relational databases into a single master table, generating a synthetic master table, and dividing it into separate databases—presents several challenges. To illustrate, consider education data as an example, with two tables (student and teacher information) linked by students’ enrollment in a teacher’s course. First, flattening often introduces numerous null values, complicating the differentiation between missing and intentionally null data. If a teacher isn’t teaching, student information in the master table will show up as null, even if the original tables had none. Second, flattening creates scalability and redundancy issues. Records in the master table often contain a high number of features, leading to longer running times for SOTA DP synthetic data generation algorithms as features increase [e.g., 32, 31, 50, 1]. Moreover, records from the original tables can be duplicated across entries, increasing redundancy and storage demand. Finally, breaking down a synthetic master table may disrupt relationships when shared feature values are involved. For instance, if the master table contains two courses with students sharing identical demographic data, it’s unclear if they’re the same student or different individuals with matching details.
Our goal is to introduce the first-of-its-kind algorithm that can generate synthetic relational databases while preserving the privacy, statistical properties, and referential integrity of the original data. For this purpose, we first discuss DP in relational databases, extend the definition of $k$-way marginal queries [46, 36] to relational databases, and articulate the requirements of referential integrity (Section 2). The main technical contribution is introducing an iterative algorithm that effectively learns the relationship between various tables (Section 3). At each iteration, the algorithm identifies a subset of $k$-way marginal queries with the highest approximation errors and then refines the relational synthetic database to minimize these errors. Notably, our algorithm avoids the need to flatten relational databases into a master table; instead, it only requires querying the relational databases to compute low-order marginal distributions, which can be done using SQL aggregate functions. Consequently, we can leverage any off-the-shelf DP mechanisms to generate synthetic data for individual tables and apply our algorithm to establish their inter-table relationship. We analyze our algorithm theoretically, establishing both DP and utility guarantees.
Existing literature on synthetic relational data generation, even without DP guarantees, remains limited [29, 34, 53, 14, 10, 54].
In summary, our contributions include:
  • Presenting a pioneering study on privacy-preserving synthetic data generation in relational databases.
  • Introducing an iterative algorithm that establishes relationships between synthetic tables while preserving statistical properties and referential integrity.
  • Developing a model-agnostic approach that integrates with any DP synthetic data algorithm, offering both privacy and utility guarantees.
@@
Privacy-preserving synthetic data generation is an active research topic [see e.g., 5, 47, 6, 11, 16, 37, 51]. For example, a line of work considered learning probabilistic graphical models [55, 30, 31] or generative adversarial networks [52, 2, 22, 44, 33, 3] with DP guarantees and generating synthetic data by sampling from these models. Another line of work proposed to iteratively refine synthetic data to minimize its approximation error on a pre-selected set of workload queries [18, 15, 32, 27, 26, 49, 1, 50, 28]. They only focused on a single-source dataset without considering relational database. The only exceptions are [53, 10]. [10] used graphical models for generating DP synthetic data, incorporating latent variables to capture the primal-foreign-key relationship. However, their method is limited to one-to-many relationships while our method can handle relationships of any type. A more closely related work is [53], which can produce multiple tables with many-to-many relationships by leveraging tools from random graph theory and representation learning. However, their method required using DP-SGD to optimize their generative models for each table generation, whereas our method can seamlessly integrate with any existing DP mechanisms for synthetic table generation. This versatility is essential, particularly as marginal-based and workload-based mechanisms often produce higher-quality synthetic tabular data than DP-SGD-based mechanisms [45, 31, 51]. Finally, some research explored DP in relational database systems, primarily focusing on releasing statistical queries [see e.g., 20, 21, 24]. In contrast, our emphasis is on releasing a synthetic copy of the relational database, extending the scope of privacy-preserving data generation methods to a more practical and general setting.
@@
Another line of work focused on the release of DP synthetic graphs [25, 56, 13, 48, 17, 4, 38, 19] with the goal of maintaining essential graph properties (e.g., connectivity, degree distribution, and cut approximation). While a relational database can often be represented as a multipartite graph (see Section 2.1), applying these existing approaches to generate synthetic relational databases encounters various challenges. First, these methods often assume that the vertex set of both real and synthetic graphs is common, learning the edge set through DP mechanisms. In our scenario, however, the vertex sets (representing records) of real and synthetic databases differ, requiring privacy preservation for both the vertex set and the edge set. Additionally, existing approaches focus on preserving graph properties, whereas our objective extends to preserve both graph properties (i.e., relationships between different tables) and statistical properties of relational databases.
@@ 2.1 Bipartite Graph and Relational Database
To simplify our presentation, we assume that the relational database $\mathcal{B}$ consists of only two tables $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ (e.g., student and teacher information), each having $n_{1}$ and $n_{2}$ rows. The relationship between tables is typically defined through primary and foreign keys. A primary key is a column in a table serving as a unique identifier for each row. A foreign key is a column that establishes a relationship between tables by referencing the primary key of a different table.
We represent the relational database as a bipartite graph $\mathcal{B}=(\mathcal{D}_{1},\mathcal{D}_{2},\bm{B})$ where each record corresponds to a node; the two tables define two distinct sets of nodes; and the relationships between the tables are depicted as edges connecting these nodes. We represent the edges using a bi-adjacency matrix, denoted as $\bm{B}\in\{0,1\}^{n_{1}\times n_{2}}$, where $B_{i,j}=1$ iff the i-th record from table $\mathcal{D}_{1}$ is connected to the j-th record from table $\mathcal{D}_{2}$ (e.g. if the i-th student is enrolled in the j-th teacher’s course). For each record $\bm{x}$, we represent its degree as the total number of records from the other table connected to $\bm{x}$, denoted by $\deg(\bm{x})$. Finally, we assume the degree of each record is upper bounded by a constant $d_{\max}$.
@@ 2.2 Differential Privacy
We first recall the definition of differential privacy (DP) [12].
Definition 1.
A randomized mechanism $\mathcal{M}$ that takes a relational database as input and returns an output from a set $\mathcal{R}$ satisfies $(\varepsilon,\delta)$-differential privacy, if for any adjacent relational databases $\mathcal{B}$ and $\mathcal{B}^{\prime}$ and all possible subsets $\mathcal{O}$ of $\mathcal{R}$, we have
$\Pr(\mathcal{M}(\mathcal{B})\in\mathcal{O})\leq e^{\varepsilon}\Pr(\mathcal{M}(\mathcal{B}^{\prime})\in\mathcal{O})+\delta.$ (1)
@@ 2.3 Problem Formulation
Our goal is to generate a privacy-preserving synthetic database $\mathcal{B}_{\text{syn}}=(\mathcal{D}_{1}^{\text{syn}},\mathcal{D}_{2}^{\text{syn}},\bm{B}^{\text{syn}})$ that preserves both statistical properties and referential integrity of the original relational database. In terms of statistical properties, our aim is to ensure that the marginal distributions of subsets of features in the synthetic data closely align with those of the real data. To achieve this, we revisit the definition of $k$-way marginal query and extend it to suit relational databases.
Definition 2.
Suppose the domain of table $\mathcal{D}_{i}$ has $d_{i}$ categorical features: $\mathcal{X}_{i}=\mathcal{X}_{i,1}\times\cdots\mathcal{X}_{i,d_{i}}$. We define a $k$-way workload as a subset of $k$ features: $\mathcal{S}\subseteq[d_{1}]$ (or $\mathcal{S}\subseteq[d_{2}]$) with $|\mathcal{S}|=k$. Additionally, given a reference value $\bm{y}$, we define a (single-table) $k$-way marginal query as
@@
where $\mathbb{I}$ is an indicator function. Analogously, we define a (cross-table) $k$-way workload as $\mathcal{S}_{1}\times\mathcal{S}_{2}\subseteq[d_{1}]\times[d_{2}]$ with $|\mathcal{S}_{1}|+|\mathcal{S}_{2}|=k$, $0<|\mathcal{S}_{1}|<k$; additionally, given a reference value $(\bm{y}_{1},\bm{y}_{2})$, we define a (cross-table) $k$-way marginal query $q_{(\mathcal{S}_{1},\mathcal{S}_{2}),(\bm{y}_{1},\bm{y}_{2})}(\bm{x}_{1},\bm{x}_{2})$ is defined as
$\prod_{(j_{1},j_{2})\in\mathcal{S}}\mathbb{I}({(x_{1,j_{1}},x_{2,j_{2}})=(y_{1,j_{1}},y_{2,j_{2}})})$
for $(\bm{x}_{1},\bm{x}_{2})\in\mathcal{X}_{1}\times\mathcal{X}_{2}$. We denote the set of all (cross-table) $k$-way marginal workloads by $\mathcal{W}_{\text{cross},k}$ and their associated queries by $\mathcal{Q}_{\text{cross},k}$.
In maintaining referential integrity, we use a linking table to establish relationships. It is designed to store the IDs of synthetic tables: the entry $(i,j)$ is included in this table if $\bm{B}_{i,j}^{\text{syn}}=1$. Moreover, when the original database exhibits a one-to-many relationship (i.e., a record in the parent table $\mathcal{D}_{1}$ can be related to one or more records in $\mathcal{D}_{2}$, but a record in the child $\mathcal{D}_{2}$ can be related to only one record in $\mathcal{D}_{1}$) or one-to-one relationship [^1] We assume that information about the types of relationships in the original database is publicly available, but it can also be learned differentially privately by examining the degree distribution of each table., we expect the synthetic database to preserve this relationship. Finally, in the case of a one-to-many relationship in the original database, we ensure there are no orphaned rows in the synthetic database. Specifically, we impose the requirement that each record in the child table $\mathcal{D}_{2}^{\text{syn}}$ connects to a record in the parent table $\mathcal{D}_{1}^{\text{syn}}$ (although records in the parent table are permitted to have no associated child records).

### v2

# Differentially Private Synthetic Data Generation for Relational Databases

@@ Abstract
Existing differentially private (DP) synthetic data generation mechanisms typically assume a single-source table. In practice, data is often distributed across multiple tables with relationships across tables. In this paper, we introduce the first-of-its-kind algorithm that can be combined with any existing DP mechanisms to generate synthetic relational databases. Our algorithm iteratively refines the relationship between individual synthetic tables to minimize their approximation errors in terms of low-order marginal distributions while maintaining referential integrity. This algorithm eliminates the need to flatten a relational database into a master table (saving space), operates efficiently (saving time), and scales effectively to high-dimensional data. We provide both DP and theoretical utility guarantees for our algorithm. Through numerical experiments on real-world datasets, we demonstrate the effectiveness of our method in preserving fidelity to the original data.
Keywords: Differential privacy, synthetic data, optimization, sampling, relational database.
@@ 1 Introduction
Relational databases play a pivotal role in modern information systems and business operations due to their efficiency in managing structured data (Schleich et al., 2019). According to a Kaggle survey (Kaggle, 2017), 65.5% of users worked extensively with relational data. Additionally, the majority of leading database management systems (e.g., MySQL and Oracle) are built on relational database principles (Ranking, 2023). These systems organize data into multiple tables, each representing a specific entity, and the relationships between tables delineate the connections between these entities. However, the widespread use of relational databases also carries a significant risk of privacy leakage. For example, if a single table suffers from a privacy breach, all other tables containing sensitive information can be exposed, as they are interconnected and share relationships. Moreover, deleting data in a relational database can be complex, and incomplete deletion can leave traces of data, which can potentially be accessed and reconstructed by attackers.
Today, differential privacy (DP) stands as the de facto standard for privacy protection (Abowd, 2018; The White House, 2023). There is a growing body of research focused on applying DP to generate private synthetic data (see, e.g., Ridgeway et al., 2021; McKenna et al., 2021; Liu et al., 2021b; SmartNoise, 2023). This effort enables data curators to share (synthetic) data while ensuring the privacy of individuals’ personal information within the original dataset. In return, they can harness advanced machine learning (ML) techniques employed by end-users trained on the synthetic data. Numerous studies have provided evidence that state-of-the-art (SOTA) DP synthetic data effectively preserves both the statistical properties of the original data and the high performance of downstream predictive models trained on these synthetic datasets when deployed on the original data (Tao et al., 2021; Wang et al., 2023). However, all existing works assume a single-source database (with a few exceptions discussed in Related Work). This motivates the central question we aim to tackle in this paper:
@@
The conventional approach—flattening relational databases into a single master table, generating a synthetic master table, and dividing it into separate databases—presents several challenges. To illustrate, consider education data as an example, with two tables (student and teacher information) linked by students’ enrollment in a teacher’s course. First, flattening may lead to data integrity issues by introducing numerous null values. This makes it difficult to distinguish between missing and intentionally null data. For instance, if a teacher is not currently teaching, their students’ information in the master table will be represented by null values. Consequently, the master synthetic table will contain a significant number of null values even if the original tables do not have any null values. Second, flattening introduces concerns related to data scalability and redundancy. Each record in the master table may have an extremely large number of features, which poses challenges for SOTA DP synthetic data generation mechanisms, as their running time grows rapidly as the number of features increases (e.g., McKenna et al., 2022, 2021; Vietri et al., 2022; Aydore et al., 2021). Additionally, certain records from the original tables might be duplicated across multiple entries in the master table, resulting in data redundancy and an increased demand for storage space. Third, breaking down a synthetic master table may disrupt relationships, especially when there are records with shared feature values. For example, if the master table includes two rows with identical student demographic information, it is unclear whether they are the same student or different students with matching demographic details. Finally, since individual records may repeat multiple times in the master table, the sensitivity of statistical queries can significantly increase, potentially becoming unbounded. Consequently, applying existing DP mechanisms may result in a synthetic master table with limited utility.
The goal of this paper is to introduce the first-of-its-kind algorithm that can generate synthetic relational databases while preserving the privacy, statistical properties, and referential integrity of the original data. For this purpose, we first extend the definition of $k$-way marginal queries (Thaler et al., 2012; Ridgeway et al., 2021) to relational databases and articulate the requirements of referential integrity (Section 2). In the context of relational databases, $k$-way marginal queries measure the low-order marginal distributions of features (i.e., columns) across different tables. Preserving these distributions is essential for many downstream applications of synthetic data, such as data visualization, building data pre-processing pipelines, and training ML predictive models like logistic regression.
Our main contribution is an algorithm that effectively generates synthetic relational data by minimizing approximation errors in terms of $k$-way marginal queries (Algorithm 1). This algorithm works as follows. First, we apply an off-the-shelf DP mechanism to generate individual synthetic tables. Then we represent the relational synthetic database as a bipartite graph and establish relationships by learning its bi-adjacency matrix. We prove that $k$-way marginals can be reformulated as (fractional) linear functions of this bi-adjacency matrix (Lemma 1). Based on this observation, we propose a combinatorial optimization that minimizes the worst-case approximation error of marginal queries between the original and synthetic databases (Section 3). The combinatorial optimization is solved using an iterative algorithm. At each step, we identify a subset of $k$-way marginal queries with the highest approximation errors. Then we solve a relaxed optimization in continuous space to update the bi-adjacency matrix (Section 4). Following this, we introduce an unbiased sampling algorithm to convert the learned bi-adjacency matrix back into discrete space (Section 5). To generate large-scale synthetic data, we propose a random slicing heuristic that leverages the sparsity of the bi-adjacency matrix, allowing us to optimize bi-adjacency matrices with millions of parameters (Section 6).
Our approach offers several benefits. First, it is highly flexible, as it can be combined with any off-the-shelf DP synthetic data generation mechanisms. These DP mechanisms are used to generate individual synthetic tables, and our algorithm can then be applied to build relationships between these tables. Second, it avoids the need to flatten relational databases into a master table; instead, it only requires querying the relational databases to compute low-order marginal distributions, which can be done using SQL aggregate functions. Third, we establish convergence guarantees for our algorithm and prove that our algorithm has linear runtime complexity with respect to the size of the synthetic data (Theorem 3 and 4). These properties, combined with the random slicing heuristic, enable running our algorithm to generate large-scale synthetic data. Finally, we present utility theorems that upper bound the approximation error for estimating $k$-way marginals using synthetic data generated by our method (Theorem 5 and 6). The proofs leverage properties of negatively associated random variables (Joag-dev and Proschan, 1983) to establish Hoeffding’s inequality for dependent variables, along with classical tools from convex geometry and concentration inequalities. We believe these new proof techniques for proving properties of synthetic data could be of independent interest to the DP synthetic data community.
We conduct numerical experiments to validate our method. We couple the iterative algorithm with SOTA DP synthetic data generation mechanisms to generate relational synthetic databases. We present a scalable PyTorch implementation of our algorithm suited for high-dimensional data. The results show that our algorithm effectively maintains key statistical properties of the original relational database to a significant extent. Note that existing literature on synthetic relational data generation, even without DP guarantees, is limited (Patki et al., 2016; Mami et al., 2022; Yang et al., 2022; Xu et al., 2023; Cai et al., 2023; Francis, 2024; Pang et al., 2024). We hope our efforts (e.g., introducing new benchmark datasets and open-sourcing PyTorch code) can benefit the communities focused on DP synthetic data and inspire new research on this topic.
@@
Privacy-preserving synthetic tabular data generation is an active research topic (see e.g., Blum et al., 2013; Ullman and Vadhan, 2020; Boedihardjo et al., 2022; Chen et al., 2015; Ge et al., 2020; Rosenblatt et al., 2020; Wang et al., 2023; He et al., 2023; Donhauser et al., 2024; Greenewald et al., 2024; Bun et al., 2024). For example, a line of work consider learning probabilistic graphical models (Zhang et al., 2017; McKenna et al., 2019, 2021) or generative adversarial networks (Xie et al., 2018; Beaulieu-Jones et al., 2019; Jordon et al., 2019; Tantipongpipat et al., 2019; Neunhoeffer et al., 2020; Bie et al., 2023) with DP guarantees and generating synthetic data by sampling from these models. Another line of work propose to iteratively refine synthetic data to minimize its approximation error on a pre-selected set of workload queries (Hardt et al., 2012; Gaboardi et al., 2014; McKenna et al., 2022; Liu et al., 2021b, a; Vietri et al., 2020; Aydore et al., 2021; Vietri et al., 2022; Liu et al., 2023b).
@@
Most existing work on DP synthetic data generation only focuses on a single-source dataset without considering relational database. The only exceptions are Xu et al. (2023); Cai et al. (2023). Among them, Cai et al. (2023) uses graphical models for generating DP synthetic data, incorporating latent variables to capture the primal-foreign-key relationship [^1] A primary key is a column in the parent table that uniquely identifies each row. A foreign key is a column in the child table that creates relationships between the two tables by referencing the primary key in the parent table.. However, their approach is restricted to one-to-many relationships (i.e., records in one table have exactly one relationship with the other table) and can only produce synthetic relational data of this type. As discussed in Appendix C, with only one-to-many relationships, the structure of relational data is simpler, making the synthetic data generation process more straightforward and computationally efficient. In contrast, our method can accommodate any type of relationship, including the more complex many-to-many cases (i.e., records in one table can be related to multiple records in another table, and vice versa). A more closely related work is Xu et al. (2023), which can produce multiple tables with many-to-many relationships by leveraging tools from random graph theory and representation learning. However, their method requires using DP-SGD to optimize their generative models for each table generation, whereas our method can seamlessly integrate with any existing DP mechanisms for synthetic table generation. This versatility is essential, particularly as marginal-based and workload-based mechanisms often produce higher-quality synthetic tabular data than DP-SGD-based mechanisms (Tao et al., 2021; McKenna et al., 2021; Wang et al., 2023). Finally, some research explores DP in relational database systems, primarily focusing on releasing statistical queries (see e.g., He, 2021; Johnson et al., 2018; Kotsogiannis et al., 2019). In contrast, our emphasis is on releasing a synthetic copy of the relational database, extending the scope of privacy-preserving data generation methods to a more practical and general setting.
@@
Another line of work focused on the release of DP synthetic graphs (Liu et al., 2023a; Zhang et al., 2021; Eliáš et al., 2020; Upadhyay, 2013; Gupta et al., 2012; Blocki et al., 2012; Sala et al., 2011; Hay et al., 2009; Liu et al., 2024) with the goal of maintaining essential graph properties (e.g., connectivity, degree distribution, and cut approximation). While a relational database can often be represented as a multipartite graph (see Section 2.1), applying these existing approaches to generate synthetic relational databases encounters various challenges. First, most existing methods assume that the vertex set does not have any attribute information and is the same between real and synthetic graphs. They only focus on learning the edge set through DP mechanisms. In our case, however, the vertex sets (representing records) include attribute information (feature values of records) and differ between real and synthetic databases, requiring privacy preservation for both the vertex attributes and the edge set. The only exceptions are Chen et al. (2019); Jorgensen et al. (2016), which consider vertex attributes when generating DP synthetic graphs. However, Chen et al. (2019) only produces synthetic graphs with the same number of vertices as the original, while applying Jorgensen et al. (2016) to generate synthetic relational databases can lead to referential integrity violation (e.g., one-to-many relationships in the original data could become many-to-many in the synthetic data). Finally, existing approaches prioritize preserving graph properties, whereas our objective extends to preserve both graph properties (i.e., relationships between different tables) and statistical properties of relational databases (i.e., low-order marginal distributions).
Broadly, a body of work has investigated applications of relational databases without imposing privacy constraints (Patki et al., 2016; Mami et al., 2022; Yang et al., 2022; Francis, 2024; Pang et al., 2024). For instance, some works focus on training predictive models on relational data (Cvitkovic, 2020; Gan et al., 2024), while others aim to analyze and extract representations from these datasets (Cai et al., 2021; Fey et al., 2023). There is also research on generating synthetic relational databases, but these efforts are typically restricted to handling one-to-many relationships defined by primary-foreign keys (Patki et al., 2016; Mami et al., 2022; Pang et al., 2024). Additionally, some studies have introduced benchmarks for evaluating deep learning models on relational databases (Robinson et al., 2024). In contrast, our work focuses on synthetic relational data generation that supports both one-to-many and many-to-many relationships. Furthermore, we incorporate DP guarantees into the data generation process to ensure robust privacy protection for the original relational databases.
@@ 2.1 Bipartite Graph and Relational Database
To simplify our presentation, we assume that the relational database $\mathcal{B}$ consists of only two tables $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ (e.g., student and teacher information), each having $n_{1}$ and $n_{2}$ rows. We represent the relational database as a bipartite graph $\mathcal{B}=(\mathcal{D}_{1},\mathcal{D}_{2},\bm{B})$ where each record corresponds to a node; the two tables define two distinct sets of nodes; and the relationships between the tables are depicted as edges connecting these nodes. We represent the edges using a bi-adjacency matrix, denoted as $\bm{B}\in\{0,1\}^{n_{1}\times n_{2}}$, where $B_{i,j}=1$ iff the $i$-th record from table $\mathcal{D}_{1}$ is connected to the $j$-th record from table $\mathcal{D}_{2}$ (e.g. if the $i$-th student is enrolled in the $j$-th teacher’s course). For each record $\bm{x}$, we represent its degree as the total number of records from the other table connected to $\bm{x}$, denoted by $\deg(\bm{x})$. Finally, we assume the degree of each record is upper bounded by a constant $d_{\max}$.
@@ 2.2 Differential Privacy
We recall the definition of differential privacy (DP) (Dwork and Roth, 2014) and provide detailed background information in Appendix A.
Definition 1
A randomized mechanism $\mathcal{M}$ that takes a relational database as input and returns an output from a set $\mathcal{R}$ satisfies $(\epsilon,\delta)$-differential privacy, if for any adjacent relational databases $\mathcal{B}$ and $\mathcal{B}^{\prime}$ and all possible subsets $\mathcal{O}$ of $\mathcal{R}$, we have
$\mathbb{P}(\mathcal{M}(\mathcal{B})\in\mathcal{O})\leq e^{\epsilon}\mathbb{P}(\mathcal{M}(\mathcal{B}^{\prime})\in\mathcal{O})+\delta.$ (1)
@@ 2.3 Problem Formulation
Our goal is to generate a privacy-preserving synthetic database $\mathcal{B}_{\text{syn}}=(\mathcal{D}_{1}^{\text{syn}},\mathcal{D}_{2}^{\text{syn}},\bm{B}^{\text{syn}})$ that preserves both statistical properties and referential integrity of the original relational database. In terms of statistical properties, our aim is to ensure that the marginal distributions of subsets of features in the synthetic data closely align with those of the real data. To achieve this, we revisit the definition of $k$-way marginal query and extend it to suit relational databases. For any integer $d\geq 1$, we define $[d]\triangleq\{1,\cdots,d\}$.
Definition 2
Suppose the domain of table $\mathcal{D}_{i}$ has $d_{i}$ categorical features: $\mathcal{X}_{i}=\mathcal{X}_{i,1}\times\cdots\mathcal{X}_{i,d_{i}}$. We define a (single-table) $k$-way workload as a subset of $k$ features: $\mathcal{S}\subseteq[d_{1}]$ (or $\mathcal{S}\subseteq[d_{2}]$) with $|\mathcal{S}|=k$. Additionally, given a reference value $\bm{y}$, we define a (single-table) $k$-way marginal query as
@@
where $\mathbb{I}$ is an indicator function. Analogously, we define a (cross-table) $k$-way workload as $\mathcal{S}_{1}\times\mathcal{S}_{2}\subseteq[d_{1}]\times[d_{2}]$ with $|\mathcal{S}_{1}|+|\mathcal{S}_{2}|=k$, $0<|\mathcal{S}_{1}|<k$; additionally, given a reference value $(\bm{y}_{1},\bm{y}_{2})$, we define a (cross-table) $k$-way marginal query as
$q_{(\mathcal{S}_{1},\mathcal{S}_{2}),(\bm{y}_{1},\bm{y}_{2})}(\bm{x}_{1},\bm{x}_{2})\triangleq\prod_{(j_{1},j_{2})\in\mathcal{S}}\mathbb{I}({(x_{1,j_{1}},x_{2,j_{2}})=(y_{1,j_{1}},y_{2,j_{2}})})\quad\text{for }(\bm{x}_{1},\bm{x}_{2})\in\mathcal{X}_{1}\times\mathcal{X}_{2}.$
We denote the set of all (cross-table) $k$-way marginal workloads by $\mathcal{W}_{\text{cross},k}$ and their associated queries by $\mathcal{Q}_{\text{cross},k}$.
Our main objective in generating synthetic relational databases is to preserve $k$-way marginal queries. These low-order statistics are useful for many downstream applications of synthetic data. For example, data visualization (e.g., histograms of each column and correlation matrices), data pre-processing pipelines (e.g., normalizing columns using their mean and standard deviation), and training simple ML models (e.g., linear models) all depend on low-order statistics. Therefore, preserving $k$-way marginals in synthetic data allows downstream users to maintain statistical insights extracted from synthetic data and train ML models that perform comparably (within a tolerance) to those trained on real data. Another example highlighting the importance of $k$-way marginals comes from the National Institute of Standards and Technology (NIST), which organized a competition in 2018 to emphasize the need for privacy preservation in synthetic data generation (McKenna et al., 2021; Ridgeway et al., 2021). In this competition, 3-way marginals were used as a utility metric for evaluating the quality of synthetic data. Here we extend the definition of $k$-way marginals to multiple synthetic tables with relationships.
In maintaining referential integrity, we use a linking table to establish relationships. It is designed to store the IDs of synthetic tables: the entry $(i,j)$ is included in this table if $B_{i,j}^{\text{syn}}=1$. Moreover, when the original database exhibits a one-to-many relationship (i.e., a record in the parent table can be related to one or more records in the child table, but a record in the child table can be related to only one record in the parent table) or one-to-one relationship [^2] We assume that information about the types of relationships in the original database (e.g., one-to-many or many-to-many) is publicly available., we expect the synthetic database to preserve this relationship. Finally, in the case of a one-to-many relationship in the original database, we ensure there are no orphaned rows in the synthetic database. Specifically, we impose the requirement that each record in the child table connects to a record in the parent table (although records in the parent table are permitted to have no associated child records).